Circular No. 7988

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002gb M. Hamuy, Carnegie Observatories, reports that a spectrum (range

380–920 nm) of SN 2002gb (cf. IAUC 7983), obtained on Oct. 8 UT with the Las Campanas Dupont 2.5-m telescope (+ WFCCD), bears resemblance to the type-Ia SN 1994D, 20 days past maximum, from which a redshift of ≈ 0.074 is derived.

A. V. Filippenko and R. Chornock, University of California, Berkeley, also write that CCD spectra (range 310–930 nm) obtained on Oct. 8 UT with the Keck-I 10-m telescope (+ Low Resolution Imaging Spectrometer) reveal that SN 2002gb is of type Ia, ∼ 1 month past maximum brightness. The redshift of the host galaxy, measured from the narrow (H II region) Hα emission line, is 0.074.

SUPERNOVAE 2001ig, 2002eg, 2002ew, 2002gc Filippenko and Chornock add that CCD spectra, obtained as above

on Oct. 8 UT, reveal that SN 2002gc (IAUC 7983) is of type Ia, but with spectral peculiarities that resemble those of SN 2002eu (IAUC 7965): the relative strengths of the Ti II/Si II 580-nm and Si II 615-nm features are similar to those of SN 1991bg, but it lacks the strong Ti II 420-nm absorption characteristic of SN 1991bg-like objects. SN 2002eg (IAUC 7944) is of type IIb, as previously reported (IAUC 7952); He I absorption lines are now easily visible. SN 2002ew (IAUC 7964) is of type II, as had been suspected, but it exhibits some spectral peculiarities, such as substantial O I 777.4-nm absorption and strong Hα emission (with essentially no absorption). SN 2001ig (IAUC 7772) has transformed itself into a type-Ib/c supernova in the nebular phase, as had been predicted (IAUC 7793). The [O I] and [Ca II] emission lines are very prominent, and Mg I] 457.1-nm is the strongest (relative to [O I] 630-nm) seen thus far in any supernova, exceeding even its strength in SN 2002ap (Leonard et al. 2002,

http://xxx.lanl.gov/abs/astro-ph/?0206368).

V4743 SAGITTARII R. J. Ivison, Royal Observatory, Edinburgh, on behalf of the James

Clerk Maxwell Telescope nova-monitoring group, reports the detection of V4743 Sgr with the SCUBA submillimeter camera. The flux density measured on Oct. 3.1 UT in good conditions was 39 ± 4 mJy at 0.85 mm, with an upper limit of 225 mJy at 0.45 mm. Very Large Array observations are planned for Oct. 11, with a flux density of ≥ 0.05 mJy expected at 3.6 cm.

2002 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green